​Real-Time Data Analysis​ Exercise*

Calculate the missing values in the table below by using the provided data collected in the household survey for

March 2019

.

​(Enter your responses as whole​ numbers.)

Category

Value

​Working-age population**

nothing

​Employed**

156 comma 748

​Unemployed**

nothing

Unemployment rate

3.8

​%

Labor​ force**

nothing

Labor force participation rate

63.0

​%

​** Numbers in thousands

​*Real-time data provided by Federal Reserve Economic Data​ (FRED), Federal Reserve Bank of Saint Louis.
Answer #1

Formulas:

Unemployment rate = Unemployed/labor force*100

Labor force = unemployed+employed

Labor force participation rate = labor force/Working age population*100

So,

3.8 = Number of unemployed/number of unemployed+156748*100

595642.4+3.8(Unemployed) = 100(Unemployed)

595642.4 = 96.2*Unemployed

Number of Unemployed = 595642.4/96.2 = 6192

Labor force = 156748+6192 = 162940

Next,

63 = 162940/Working age population*100

Working age population = 162940/63*100

=258635